OpenStudy (anonymous):

two cars leave Indianapolis, one traveling east and the other west. After 3 hours they are 297 miles apart. If one car is traveling 5 mph faster than the other, what is the speed of each?

OpenStudy (anonymous):

52 and 47

OpenStudy (anonymous):

d= rt, Let x = speed of first car First car d1 = x * 3 Second car d2 = (x - 5)* 3 d1 + d2 = 297 so, 3x + 3 (x-5) = 297 x = 52 another car x - 5 = 52-5 = 47 Answer is 47, 52mph